OpenStudy (anonymous):

Write a system of two equations using x and y in each for the following. The sum of two numbers is 170 and their difference is 26.

OpenStudy (anonymous):

x+y=170 and y-x=26 y-x=26 y=26+x x+26+x=170 2x=170-26 2x=144 x=72 x+y=170 72+y=170 y=170-72 y=98

OpenStudy (anonymous):

1. x+y = 170 2. x-y = 26